OpenStudy (anonymous):

An object's acceleration is given by the formula a(t) = 0.8 + 0.4e^0.2t where a is the acceleration of the object in ms-2 and t is the time in seconds since the start of the objects motion. if the object had a velocity of 4ms-1 after 3 seconds, how far did it travel between t = 5 and t = 6?

OpenStudy (anonymous):

I got v = 0.8t + 2e^0.2t + c when i integrate the formula

OpenStudy (primeralph):

Use the boundary condition given.

OpenStudy (primeralph):

Plug in 3 for t and 4 for v. Find c.

OpenStudy (anonymous):

v(3) = 4 2.4+2e^0.6+c = 4 c = 1.6-2e^0.6 c=-2.04 is that right?

OpenStudy (primeralph):

If you did it right, then yes. Now find the average distance traveled.

OpenStudy (anonymous):

Is it alright if you double check for me?

OpenStudy (anonymous):

just in case I made a mistake that i cant see

OpenStudy (anonymous):

average distance i got was 8.38

OpenStudy (primeralph):

(V(6)-V(5))* (6-5)